If A and B are independent events with P(A)=0.90 and P(A AND B)=0.54, find P(B). Give your answer as a decimal rounded to two decimal places.
Answer #1

Since A and B are independent events,

Then

P( A AND B) = P(A) P (B)

0.54=(0.9) P(B)

P(B) = 0.54/0.9

= 54/90

= 9/15

= 3/5

= 0.60

P(B)= 0.60.